# Source Code Foundation Models are Transferable Binary Analysis Knowledge Bases

### v1

@@ 1 Introduction
In recent years, we see two trends of uni-modal code model pre-training. On one hand, there is a remarkable surge in the development of generative Source Code Foundation Models (SCFMs) [12, 64, 55, 22, 43], along with advancements in general Large Language Models (LLMs) [4, 59, 46]. Driven by a growing interest in automating software development, these powerful models are trained on billions of tokens from diverse codebases, covering a wide spectrum of programming languages [35]. They possess the capability to complete, infill [20], and refine code [66], as well as generate code from natural language instructions [44]. On the other hand, there is a stream of research focusing on binary understanding models [60, 57], which target learning nuanced code semantics with structures of low-level code, which is critical for software security. Both fields are evolving through continuous pre-training on expansive datasets of uni-modal data, setting new benchmarks in both effectiveness and complexity.
@@
Human-Oriented Binary Reverse Engineering (HOBRE), which involves automatically lifting binary to human understandable contents [14, 68], typically source-code related, occupies a unique intersection between these two fields. Existing decompilers for binary reverse engineering are able to translate binary code to C-style code that is functionally equivalent to its original source code. However, a significant semantic gap remains between the decompiled code and its original source, primarily due to the absence of meaningful symbolic information in binary code. Hence, human expertise is still indispensable in the reverse engineering process. HOBRE aims to bridge this semantic gap, which traditionally requires substantial human effort, by leveraging cross-modal deep learning models. Existing approaches either train task-specific small expert models in a supervised manner [32, 1, 68], which lack generalizability as shown in later evaluations [56], or require extensive continual pre-training of uni-modal SCFMs [30] which is undesirable considering cost and the risk of forgetting previously acquired source code knowledge [34, 72]. There are also attempts in directly prompting LLMs for HOBRE, which, even though demonstrates better generalizability than small supervised models, also face challenges in understanding stripped decompiled code that lacks symbolic information [31].
@@
Our insight is that this semantic gap between binary and source code is analogous to the gap between low-level pixels in images and high-level concepts in natural language, which can be bridged with sufficient understanding of both. Inspired by the achievements of multi-modal models that seamlessly integrate vision, audio, or other signals with language to facilitate reasoning [2, 39, 42, 45], we hypothesize that HOBRE could similarly benefit from leveraging uni-modal models developed for both source code and binary code. Such integration would enhance our ability to bridge the semantic gap and enable more effective semantic lifting.
@@
In this paper, we validate this idea by proposing a novel probe-and-recover framework ProRec that incorporates a binary-source encoder-decoder model and black-box LLMs for HOBRE, featuring a compute-efficient cross-modal alignment approach of a binary function encoder and a frozen SCFM for the binary-source model. The workflow of ProRec is shown in Figure 1. The aligned binary-source model acts as a cross-modal-knowledge-prober that can synthesize symbol-rich, diverse source code fragments condition on binary input, denoted as probed contexts. The black-box LLM functions as recoverer that takes as input the binary function together with the probed contexts for tasks such as binary summarization. Intuitively, the conditional source code synthesis by the aligned binary-source code model can be viewed as probing the base SCFM as a parametric knowledge base [52] with a binary function as query, given that the SCFM’s weights remains unchanged before and after the alignment. A black-box LLM analyzes and aggregates these knowledgable contexts with the binary function for recovery. This way, ProRec leverages both cross-modal aligned knowledge and strong reasoning ability of LLMs and can outperform directly letting the LLM to reason. ProRec is general and can be applied to different base architectures, continually evolve with base models.
We demonstrate the effectiveness of ProRec on two core tasks in reverse engineering [9, 10]: binary summarization and binary function name recovery. The former aims to generate natural language descriptions for a binary function, and the later aims to recover the function name of a decompiled function. We evaluate ProRec on a diversified dataset compiled from GitHub repositories, demonstrating improvements of 3.1% (10.3% relative gain) in CHRF and 12% (16.7% relative gain) in a GPT4-based metric that has high correlation with human judgement on the summarization task over zero-shot baseline. We conduct human study to show the effectiveness of the newly proposed GPT4-based metric. On name recovery tasks, ProRec significantly improves over zero-shot baseline by 6.7% and 7.4% for token-level precision and recall, respectively. For both tasks, ProRec also consistently show advantage over a retrieval-augmented baseline with a strong cross-modal dense retriever.  [^1] Our code and data are available at https://github.com/ziansu/prorec.
@@ 2.1 The Probe-and-Recover Framework
The decomposition is similar to that of retrieval-augmented generation [38, 67], where $p(y|x)=\sum_{s\in\text{top-}k(S^{*})}P(y|s,x)P(s|x)$, given a document pool $S^{*}$. However, there are two major differences. First, the source code fragments $\mathcal{S}_{k}$ are not retrieved from $S^{*}$, instead, they are sampled from the conditional distribution of the prober $P_{\theta,\psi}(\cdot|x)$. Due to the alignment strategy (discussed in §2.3), source code fragments sampled from the prober’s distribution have more flexibility than those retrieved a fixed document pool in binary reverse engineering scenario, potentially less noisy. We empirically demonstrate the superiority of probing over retrieval for augmentation in §4.
Second, we stress the internal analysis from the LLM denoted as $P_{\phi}(\mathcal{A}|\mathcal{S}_{k},x)$ in the decomposition. The insight is that, even though high-level recovery requires additional domain information to hint black-box LLMs for further induction, that doesn’t necessarily mean LLMs totally lacks of such knowledge (since proprietary LLMs can be significantly larger than open-source code language models in size and may have more training data, just different mixtures), it might be some long-tail knowledge that requires better prompting to exploit [37, 71]. On the other hand, the analysis help LLMs to be less influenced by the noisy contexts. This is beneficial for both retrieved and probed contexts.
@@
Figure 3: Average negative log-likelihood of source function tokens estimated by base SCLM and that of source function conditioned on its binary counterpart estimated by the aligned prober.
@@ 2.2 Model Architecture and Instantiation
The core of ProRec is the cross-modal prober, which is an encoder-decoder model aligned in the token embedding space of the SCFM, as illustrated in Figure 2. We would like to benefit from both pre-trained binary function encoders that possesses binary domain knowledge and the strong SCFMs for generalizable probing. We choose the state-of-the-art CodeArt [57] as our structure-aware binary function encoder $g(\cdot)$. CodeArt is a BERT-like transformer encoder that takes as input a disassembled binary function $x_{\texttt{asm}}$ along with its dependency graph $G_{\texttt{dep}}$ obtained by program analysis, and outputs the embeddings for all assembly code tokens and graph node tokens (each graph node token corresponds to one instruction, e.g., mov rax, [rbp+var_40], in the assembly code). We choose the Code-Llama [55] family as our base SCFM  [^3] We also tried other SCFMs like DeepSeek-Coder [22] or StarCoder2 [43] in our preliminary study and find that Code-Llama performs best as a base SCFM for our prober. .
@@
We leverage proprietary black-box LLMs (GPT3.5, Claude-3, and Gemini-1.0) as our recoverer, since they have strong reasoning ability and support long contexts. Specifically, the LLMs are prompted with $x_{\texttt{dec}}$ as zero-shot baseline. For retrieval-augmented baseline and ProRec, we append the additional context to the original input and instruct LLMs to analyze relevance and then generate recovery. Detailed prompts can be found in Appendix C.
@@ 2.3 Prober Training
Since CodeArt is exclusively pre-trained on binary code corpus [57], we first align it with a pre-trained source code encoder codet5p-embedding-110m [63] in the function-level embedding space as a pre-alignment stage in order to facilitate the later encoder-decoder alignment. To achieve this, we add a projection head for each encoder to project their [CLS] token embeddings to the same dimension $d_{\text{enc}}$, forming a standard dual-encoder [29]. This dual-encoder can encode $(x_{\texttt{asm}},G_{\text{dep}})$ into $\boldsymbol{h}_{\texttt{asm}}\in\mathbb{R}^{d_{\text{enc}}}$ and $x_{\texttt{src}}$ into $\boldsymbol{h}_{\texttt{src}}\in\mathbb{R}^{d_{\text{enc}}}$. We train the dual-encoder in a CLIP-like symmetric contrastive fashion [54]. Since the implementation is relatively standard, we refer readers to Appendix A for details.
The dual-encoder can function as a dense retriever to rank and retrieve the top-$k$ source functions from the source function pool of the training set for a query binary function, based on the similarity measure $\text{sim}(x_{\texttt{asm}},x_{\texttt{src}})=\cos\left(\boldsymbol{h}_{\texttt{asm}},\boldsymbol{h}_{\texttt{src}}\right)$. It achieves 84% recall@1 on validation set with a pool of 10k examples, demonstrating strong performance as a retriever. We utilize this dual-encoder to set up a retrieval-augmented recovery baseline to compare with ProRec in §4.
@@
The limited amount of trainable parameters results in efficient training. For memory efficiency, we apply quantization (4bit or 8bit) [17, 18] to the base SCFM during alignment.
One evidence that the knowledge of the aligned prober is mainly from the SCFM pre-training instead of learned during alignment is shown in Figure 3. We sampled 500 $(x_{\texttt{asm}},x_{\texttt{src}})$ pairs from the validation set and find that the average negative log-likelihood $-\frac{1}{|x_{\texttt{src}}|}\log P_{\psi}(x_{\texttt{src}})$ for $x_{\texttt{src}}$ provided by the base SCFM and $-\frac{1}{|x_{\texttt{src}}|}\log P_{\theta,\psi}(x_{\texttt{src}}|x_{\texttt{asm}},G_{\text{dep}})$ for $x_{\texttt{src}}$ conditioned on the $x_{\texttt{asm}}$ provided by the aligned prober are highly correlated, indicating that the prober’s ability is consistent with the base SCFM. Another interesting observation is that, instruction-tuned SCFMs typically show higher losses during alignment than their original models, which also implies the significance of pre-trained knowledge of source code for cross-modal ability as instruction-tuning may cause forgetting.
@@ 2.4 Cross-Modal Knowledge Probing
For the probing process, i.e., sampling $\mathcal{S}_{k}$ with the aligned $P_{\theta,\psi}(\cdot|x_{\texttt{asm}},G_{\text{dep}})$, we want to cover a diverse yet relevant set of candidates. We leverage nucleus sampling [26] to first let the prober generate a relatively large set of source function signatures with high randomness (top-$p=0.75$). We use idea similar to retrieval by training a binary-signature dual-encoder to rank generated signatures and filter out the noisy ones. Ultimately, we use the prober to further complete the remaining signatures with smaller randomness (top-$p=0.5$). Since signature is short and important for HOBRE, our strategy achieves both better relevance compared to using a fixed small $p$ for full function generation and better efficiency compared to sampling a large set of functions with a large $p$.
@@ 3 Experiment Setup
We evaluate ProRec on two binary reverse engineering tasks: recovering function names from decompiled code (§3.2), and summarizing function semantics from decompiled code (§3.1). In this section, we first introduce our dataset, and the setups of each task.
@@
The training and evaluation of ProRec requires pair-wise data between a binary function and its corresponding source code. To the best of our knowledge, there is no publicly available dataset that contains matched source code with the binary program. Therefore, we follow a widely adapted practice in the reverse engineering domain [13, 36, 7], using GHCC  [^4] https://github.com/huzecong/ghcc to automatically clone and compile repositories from GitHub. After the compilation, we map the resulting binary programs with the corresponding source code functions leveraging the debug information in binary programs. In total, our data consists of 270k pairs of binary and source code functions. We split 260k data samples for training and 10k data samples for test. We use 5% of the training data as the validation dataset. To make the evaluation cost tractable, we randomly sample 1k samples from the test dataset. For details in data processing and quality assurance, please see Appendix B.1.
@@ 3.1 Binary Summarization
A binary summarization tool takes as input a snippet of decompiled code, and outputs natural language descriptions. It facilitates two key processes in the reverse engineering practice [9]: understanding the purpose of a program, and understanding the functionality of a program. Please see Appendix E for a detailed example.
@@
We use the summary for a source code function as the reference summary for the corresponding decompiled code, following existing work [56]. We instruct an LLM to summarize decompiled code with three setups: (1) providing the model with only the decompiled code; (2) additionally provide the relevant source code snippets retrieved by a retriever; (3) additionally provide and the source code snippets generated by ProRec. The first two setups are considered baseline approaches for comparison.
@@
We use CHRF [53] instead of other commonly used metrics in summarization such as BLEU [48]. Our user study in Appendix D shows that only CHRF is consistent with human preferences. We additionally propose GPT4-based metrics for binary summarization, since recent studies show that LLM-based metrics show higher consistency with human judgement [73]. We ask GPT4 to measure the context and functionality relevance of a summary. The questions are derived from a thorough study on human reverse engineers [9]. Details are in Appendix B.2. The proposed metrics show high correlation with human judgement. Details are in Appendix D.

### v2

@@ 1 Introduction
In recent years, we see two trends of uni-modal code model pre-training. On one hand, there is a remarkable surge in the development of generative Source Code Foundation Models (SCFMs) Chen et al. (2021); Wang et al. (2021); Rozière et al. (2024); Guo et al. (2024); Lozhkov et al. (2024), along with advancements in general Large Language Models (LLMs) Anil et al. (2023); Touvron et al. (2023); OpenAI (2024). Driven by a growing interest in automating software development, these powerful models are trained on billions of tokens from diverse codebases, covering a wide spectrum of programming languages Kocetkov et al. (2022). They possess the capability to complete, infill Fried et al. (2022), and refine code Wei et al. (2023), as well as generate code from natural language instructions Luo et al. (2023). On the other hand, there is a stream of research focusing on binary understanding models Wang et al. (2022a); Su et al. (2024), which target learning nuanced code semantics with structures of low-level code, which is critical for software security. Both fields are evolving through continuous pre-training on expansive datasets of uni-modal data, setting new benchmarks in both effectiveness and complexity.
@@
Human-Oriented Binary Reverse Engineering (HOBRE), which involves automatically lifting binary to human understandable contents Chen et al. (2022b); Xiong et al. (2023), typically source-code related, occupies a unique intersection between these two fields. Existing decompilers for binary reverse engineering are able to translate binary code to C-style code that is functionally equivalent to its original source code. However, a significant semantic gap remains between the decompiled code and its original source, primarily due to the absence of meaningful symbolic information in binary code. Hence, human expertise is still indispensable in the reverse engineering process. HOBRE aims to bridge this semantic gap, which traditionally requires substantial human effort, by leveraging cross-modal deep learning models. Existing approaches either train task-specific small expert models in a supervised manner Jin et al. (2022); Al-Kaswan et al. (2023); Xiong et al. (2023); Ye et al. (2023b), which lack generalizability as shown in later evaluations Shang et al. (2024), or require extensive continual pre-training of uni-modal SCFMs Jiang et al. (2023) which is undesirable considering cost and the risk of forgetting previously acquired source code knowledge Kirkpatrick et al. (2017); Zhai et al. (2023). There are also attempts in directly prompting LLMs for HOBRE, which, even though demonstrates better generalizability than small supervised models, also face challenges in understanding stripped decompiled code that lacks symbolic information Jin et al. (2023).
@@
Our insight is that this semantic gap between binary and source code is analogous to the gap between low-level pixels in images and high-level concepts in natural language, which can be bridged with sufficient understanding of both. Inspired by the achievements of multi-modal models that seamlessly integrate vision, audio, or other signals with language to facilitate reasoning Alayrac et al. (2022); Li et al. (2023); Liu et al. (2024); Moon et al. (2023), we hypothesize that HOBRE could similarly benefit from leveraging uni-modal models developed for both source code and binary code. Such integration would enhance our ability to bridge the semantic gap and enable more effective semantic lifting.
@@
In this paper, we validate this idea by proposing a novel probe-and-recover framework ProRec that incorporates a binary-source encoder-decoder model and black-box LLMs for HOBRE, featuring a compute-efficient cross-modal alignment approach of a binary function encoder and a frozen SCFM for the binary-source model. The workflow of ProRec is shown in Figure 1. The aligned binary-source model acts as a cross-modal-knowledge-prober that can synthesize symbol-rich, diverse source code fragments condition on binary input, denoted as probed contexts. The black-box LLM functions as recoverer that takes as input the binary function together with the probed contexts for tasks such as binary summarization. Intuitively, the conditional source code synthesis by the aligned binary-source code model can be viewed as probing the base SCFM as a parametric knowledge base Petroni et al. (2019) with a binary function as query, given that the SCFM’s weights remains unchanged before and after the alignment. A black-box LLM analyzes and aggregates these knowledgable contexts with the binary function for recovery. This way, ProRec leverages both cross-modal aligned knowledge and strong reasoning ability of LLMs and can outperform directly letting the LLM to reason. ProRec is general and can be applied to different base architectures, continually evolve with base models.
We demonstrate the effectiveness of ProRec on two core tasks in reverse engineering Bryant (2012); Burk et al. (2022): binary summarization and binary function name recovery. The former aims to generate natural language descriptions for a binary function, and the later aims to recover the function name of a decompiled function. We evaluate ProRec on a diversified dataset compiled from GitHub repositories, demonstrating improvements of 3.1% (10.3% relative gain) in CHRF and 12% (16.7% relative gain) in a GPT4-based metric that has high correlation with human judgement on the summarization task over zero-shot baseline. We conduct human study to show the effectiveness of the newly proposed GPT4-based metric. On name recovery tasks, ProRec significantly improves over zero-shot baseline by 6.7% and 7.4% for token-level precision and recall, respectively. For both tasks, ProRec also consistently show advantage over a retrieval-augmented baseline with a strong cross-modal dense retriever.  [^1] Our code and data are available at https://github.com/ziansu/prorec.
@@ 2.1 The Probe-and-Recover Framework
The decomposition is similar to that of retrieval-augmented generation Lewis et al. (2020); Weijia et al. (2023), where $p(y|x)=\sum_{s\in\text{top-}k(S^{*})}P(y|s,x)P(s|x)$, given a document pool $S^{*}$. However, there are two major differences. First, the source code fragments $\mathcal{S}_{k}$ are not retrieved from $S^{*}$, instead, they are sampled from the conditional distribution of the prober $P_{\theta,\psi}(\cdot|x)$. Due to the alignment strategy (discussed in §2.3), source code fragments sampled from the prober’s distribution have more flexibility than those retrieved a fixed document pool in binary reverse engineering scenario, potentially less noisy. We empirically demonstrate the superiority of probing over retrieval for augmentation in §4.
Second, we stress the internal analysis from the LLM denoted as $P_{\phi}(\mathcal{A}|\mathcal{S}_{k},x)$ in the decomposition. The insight is that, even though high-level recovery requires additional domain information to hint black-box LLMs for further induction, that doesn’t necessarily mean LLMs totally lacks of such knowledge (since proprietary LLMs can be significantly larger than open-source code language models in size and may have more training data, just different mixtures), it might be some long-tail knowledge that requires better prompting to exploit Levine et al. (2022); Yu et al. (2022). On the other hand, the analysis help LLMs to be less influenced by the noisy contexts. This is beneficial for both retrieved and probed contexts.
@@
Figure 3: Negative log-likelihoods of source functions estimated by base SCLM and those conditioned on its binary counterpart estimated by the aligned prober.
@@ 2.2 Model Architecture and Instantiation
The core of ProRec is the cross-modal prober, which is an encoder-decoder model aligned in the token embedding space of the SCFM, as illustrated in Figure 2. We would like to benefit from both pretrained binary function encoders that possesses binary domain knowledge and the strong SCFMs for generalizable probing. We choose the state-of-the-art CodeArt Su et al. (2024) as our structure-aware binary function encoder $g(\cdot)$. CodeArt is a BERT-like transformer enccoder that takes as input a disassembled binary function $x_{\texttt{asm}}$ along with its dependency graph $G_{\texttt{dep}}$ obtained by program analysis, and outputs the embeddings for all assembly code tokens and graph node tokens (each graph node token corresponds to one instruction, e.g., mov rax, [rbp+var_40], in the assembly code). We choose the Code-Llama Rozière et al. (2024) family as our base SCFM  [^3] We also tried other SCFMs like DeepSeek-Coder Guo et al. (2024) or StarCoder2 Lozhkov et al. (2024) in our preliminary study and find that Code-Llama performs best as a base SCFM for our prober. .
@@
We leverage proprietary black-box LLMs (GPT3.5, Claude-3, and Gemini-Pro) as our recoverer, since they have strong reasoning ability and support long contexts. Specifically, the LLMs are prompted with $x_{\texttt{dec}}$ as zero-shot baseline. For retrieval-augmented baseline and ProRec, we append the additional context to the original input and instruct LLMs to analyze relevance and then generate recovery. Detailed prompts can be found in Appendix D.
@@ 2.3 Prober Training
Since CodeArt is exclusively pre-trained on binary code corpus Su et al. (2024), we first align it with a pre-trained source code encoder codet5p-embedding-110m Wang et al. (2023) in the function-level embedding space as a pre-alignment stage in order to facilitate the later encoder-decoder alignment. To achieve this, we add a projection head for each encoder to project their [CLS] token embeddings to the same dimension $d_{\text{enc}}$, forming a standard dual-encoder Jiang et al. (2024). This dual-encoder can encode $(x_{\texttt{asm}},G_{\text{dep}})$ into $\boldsymbol{h}_{\texttt{asm}}\in\mathbb{R}^{d_{\text{enc}}}$ and $x_{\texttt{src}}$ into $\boldsymbol{h}_{\texttt{src}}\in\mathbb{R}^{d_{\text{enc}}}$. We train the dual-encoder in a CLIP-like symmetric contrastive fashion Radford et al. (2021). Since the implementation is relatively standard, we refer readers to Appendix A for details.
The dual-encoder can function as a dense retriever to score and retrieve the top-$k$ source functions for a query binary function from the source function pool of the training set, based on the similarity measure $\text{sim}(x_{\texttt{asm}},x_{\texttt{src}})=\cos\left(\boldsymbol{h}_{\texttt{asm}},\boldsymbol{h}_{\texttt{src}}\right)$. It achieves 84% recall@1 on the validation set with a pool of 10k examples, demonstrating strong performance as a retriever. We utilize this dual-encoder to set up a retrieval-augmented HOBRE baseline to compare with ProRec in §4.
@@
The limited amount of trainable parameters results in efficient training. For memory efficiency, we apply quantization (4bit or 8bit) Dettmers et al. (2022, 2024) to the base SCFM during alignment.
One evidence that the knowledge of the aligned prober is mainly from the SCFM pre-training instead of learned during alignment is shown in Figure 3. We sampled 500 $(x_{\texttt{asm}},x_{\texttt{src}})$ pairs from the validation set and find that the negative log-likelihood $-\log P_{\psi}(x_{\texttt{src}})$ for $x_{\texttt{src}}$ provided by the base SCFM and $-\log P_{\theta,\psi}(x_{\texttt{src}}|x_{\texttt{asm}},G_{\text{dep}})$ for $x_{\texttt{src}}$ conditioned on the $x_{\texttt{asm}}$ provided by the aligned prober are highly correlated, indicating that the prober’s ability is consistent with the base SCFM. Another interesting observation is that, instruction-tuned SCFMs typically show higher losses during alignment than their original models, which also implies the significance of pre-trained knowledge of source code for cross-modal ability as instruction-tuning may cause forgetting.
@@ 2.4 Cross-Modal Knowledge Probing
For the probing process, i.e., sampling $\mathcal{S}_{k}$ with the aligned $P_{\theta,\psi}(\cdot|x_{\texttt{asm}},G_{\text{dep}})$, we want to cover a diverse yet relevant set of candidates. We leverage nucleus sampling Holtzman et al. (2019) to first let the prober generate a relatively large set of source function signatures with high randonmess (top-$p=0.75$). We use idea similar to retrieval by training a binary-signature dual-encoder to rank generated signatures and filter out the noisy ones. Ultimately, we use the prober to further complete the remaining signatures with smaller randomness (top-$p=0.5$). Since signature is short and important for HOBRE, our strategy achieves both better relevance compared to using a fixed small $p$ for full function generation and better efficiency compared to sampling a large set of functions with a large $p$.
@@ 3 Experiment Setup
We evaluate ProRec on two binary reverse engineering tasks: summarizing function semantics from decompiled code (§3.1), and recovering function names from decompiled code (§3.2). In this section, we first introduce our dataset, and the setups of each task.
@@
The training and evaluation of ProRec requires pair-wise data between a binary function and its corresponding source code. To the best of our knowledge, there is no publicly available dataset that contains matched source code with the binary program. Therefore, we follow a widely adapted practice in the reverse engineering domain Chen et al. (2022a); Lacomis et al. (2019); Banerjee et al. (2021), using GHCC  [^4] https://github.com/huzecong/ghcc to automatically clone and compile repositories from GitHub. After the compilation, we map the resulting binary programs with the corresponding source code functions leveraging the debug information in binary programs. In total, our data consists of 270k pairs of binary and source code functions. We split 260k data samples for training and 10k data samples for test. We use 5% of the training data as the validation dataset. To make the evaluation cost tractable, we randomly sample 1k samples from the test dataset. For details in data processing and quality assurance, please see Appendix B.1.
@@ 3.1 Binary Summarization
A binary summarization tool takes as input a snippet of decompiled code, and outputs natural language descriptions. It facilitates two key processes in the reverse engineering practice Bryant (2012): understanding the purpose and domain of a program (referred to as context relevance), and understanding the functionality of a program (functionality). Please see Appendix F for a detailed example.
@@
We instruct an LLM to summarize decompiled code with three setups: (1) providing the model with only the decompiled code; (2) additionally providing the relevant source code snippets retrieved from the datastore consisting of all source functions in prober’s training set by the cross-modal retriever; (3) additionally providing and the source code snippets generated by ProRec. The first two setups are considered baseline approaches for comparison. We further instruct each LLM to summarize the source code corresponding to the test samples as reference summaries.
@@
Given the reverse engineering nature of binary summarization, the automatic evaluation metrics should reflect context relevance and functionality of the summary, different from text summarization. For final results, we report CHRF Popović (2015), which our meta-evaluation (described next) identified as the most aligned with human preferences among popular existing metrics such as BLEU Papineni et al. (2002). Additionally, we introduce and report two GPT4-based metrics for context relevance and functionality judgement respectively, following LLM as a Judge Zheng et al. (2024), which demonstrate strong correlation with human judgments. The GPT4-based metrics range from 1 (worst) to 5 (best) based on corresponding criteria. Further details (e.g., prompts and rationale) about the GPT4-based metrics can be found in Appendix B.2 and Appendix D.
User Study
We conduct a user study  [^5] We obtained an IRB for the study (IRB-2024-799). to gather human judgments on the quality of binary summarization, which serves as the gold standard for this task. The study aims to (1) perform a meta-evaluation of automatic metrics and (2) accurately assess the performance of different summarization approaches. Participants are asked to score a summary based on decompiled code, corresponding source code, and the reference summary. The scoring is done on two criteria—context relevance and functionality—on a scale from 1 (worst) to 5 (best). The method used to generate each summary is not disclosed to the participants. For the meta-evaluation of automatic metrics, we calculate the Spearman correlation between human judgments and automatic metric scores. For more details, we refer readers to Appendix E.